Problem 15-10
Optimal Capital Structure with Hamada

Beckman Engineering and Associates (BEA) is considering a change in its capital structure. BEA currently has $20 million in debt carrying a rate of 8%, and its stock price is $40 per share with 2 million shares outstanding. BEA is a zero growth firm and pays out all of its earnings as dividends. The firm's EBIT is $12.168 million, and it faces a 30% federal-plus-state tax rate. The market risk premium is 5%, and the risk-free rate is 7%. BEA is considering increasing its debt level to a capital structure with 40% debt, based on market values, and repurchasing shares with the extra money that it borrows. BEA will have to retire the old debt in order to issue new debt, and the rate on the new debt will be 11%. BEA has a beta of 1.1.

  1. What is BEA's unlevered beta before restructuring? Use market value D/S (which is the same as wd/ws) when unlevering. Round your answer to two decimal places.

  2. What are BEA's new beta after releveraging and cost of equity if it has 40% debt? Round your answers to two decimal places.
    Beta
    Cost of equity %

  3. What is BEA's WACC after releveraging? Round your answer to two decimal places.
    %

    What is the total value of the firm with 40 % debt? Enter your answer in millions. For example, an answer of $1.2 million should be entered as 1.2, not 1,200,000. Round your answer to three decimal places.
    $   million
Answer #1

(9) in un levered E Bu = Bex AL D(1-t) +E the given Question BL = 11 E = 2x40 = 80 million D = 20 million t = 30% 0.8 So Bu=1situation debt in to this you now here portion raises So E = 0.6 D = 0.4 So B = 0.94 x 0.4(1-0.3) +0.6 - 1.38 I 0.6 1.38 CostSo WACC = 0.468 (1-0.3 +0.6 x 13.9. = 2.24 + 8:34 = 10.58% firm: Total value of EBIT = $12.168 million Debt portion = you Tax